OpenStudy (anonymous):

When a mortar shell is fired with an initial velocity of v0 ft/sec at an angle α above the horizontal, then its position after t seconds is given by the parametric equations x = (v0 cos α)t , y = (v0 sin α)t − 16t^2. If the mortar shell hits the ground 2500 feet from the mortar when α = 75degress, determine v0 *answer choices* 1. v0 = 360 ft/sec 2. v0 = 370 ft/sec 3. v0 = 380 ft/sec 4. v0 = 390 ft/sec 5. v0 = 400 ft/sec

OpenStudy (dumbcow):

from x equation: \[t = \frac{x}{v*\cos \theta} = \frac{2500}{v*\cos 75}\] substitute this into y equation \[0 = v*\sin 75(\frac{2500}{v*\cos 75}) -16(\frac{2500}{v*\cos 75})^{2}\] solve for v v = 400